OpenStudy (anonymous):

the perimeter of a semicircle is 4pi+8. the area of a square , the center of which lies in line with the center of the semicircle is?

OpenStudy (anonymous):

ans is 16...but dont know how to get it!

OpenStudy (anonymous):

perimeter of semi-circle=\[\pi r+d\] hence,\[\pi r+d=4\pi+8\] comparing both sides r=4 & d=8 so length of diagonal of square= 8 hence area of square= 8*8/2=32

OpenStudy (anonymous):

i dont know what is my mistake in calculating the area but the length of diagonal is 8 only

OpenStudy (anonymous):

i solved it the same way....but apparently anwer is 16

OpenStudy (anonymous):

ya, length of daimeter is 8 only

OpenStudy (anonymous):

but the it says center of square inline with center of semicircle....this is not making sense

OpenStudy (anonymous):

if we consider the line is joining the 2 sides of square then also the side of square is 8 and area would be 8*8=64